For each of the following annuities, calculate the annuity payment.

Annuity Payment Future Value Years Interest Rate
$24,450 8 8%
$990,000 39 10%
$808,000 25 11%
$134,000 16 7%
Answer #1
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
24450= Cash Flow*(((1+ 8/100)^8-1)/(8/100))
Cash Flow = 2298.66
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
990000= Cash Flow*(((1+ 10/100)^39-1)/(10/100))
Cash Flow = 2466.07
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
808000= Cash Flow*(((1+ 11/100)^25-1)/(11/100))
Cash Flow = 7062.12
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
134000= Cash Flow*(((1+ 7/100)^16-1)/(7/100))
Cash Flow = 4804.92